The acceleration of a particle is given by a relation a=v3 . It is known that at time t=0, position is -2m and velocity is CRN. Find the displacement, position, velocity and acceleration at instant of CRN/5 sec


1
Expert's answer
2021-06-29T08:17:52-0400

"x=2\cdot 5=10~m,"

"\Delta x=x_2-x_1=10-(-2)=12~m,"

"v=\frac{\Delta x}{t}=\frac{12}{5}=2.4~m,"

"a=\frac vt=\frac{2.4}{5}=0.48~\frac{m}{s^2}."